The average age of my family consisting of parents and myself is 30 year. Ten year ago, the average age of my parents were 23year.Find my present age.
Answer by JulietG(1812)

Let A be your age:
A + B + C = 90 (3x30, the average age; there are 3 of you)
(B-10)+(C-10)= 46 (23x2 people), Add 20 to each side, B+C=66
Substitute the known value of B+C from the second equation into the first equation.
A + 66 = 90
Subtract
A = 90- 66 = 24
.
Check my math; I haven't had my coffee yet ;-)